Circular No. 8791

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2007 A1 (LOVAS) Comet P/1986 W1 (= 1986 XIII = 1986p; cf. IAUC 4276) has been

accidentally rediscovered on Catalina Sky Survey images taken with the 0.68-m Schmidt telescope on Jan. 9 by R. A. Kowalski, the comet appearing diffuse with a slightly condensed core ≈ 1′ in diameter, and a more intense core ≈ 20′′ in diameter. Confirming CCD frames in good seeing by E. J. Christensen with the Mt. Lemmon 1.5-m reflector on the same night show a moderately condensed 10′′ coma with a diffuse tail extending 30′′ toward p.a. ∼ 75^{o}, while his images on Jan. 10 display a moderately condensed 10′′ inner coma surrounded by a faint 30′′ outer coma (the latter being elongated toward p.a. 60^{o}). Following posting on the Minor Planet Center’s ‘NEOCP’ website, additional astrometry below was reported by G. Hug, B. Twarog, and D. Tibbets (Eskridge, KS, 0.7-m reflector; object diffuse) and by J. E. McGaha (Tucson, AZ, 0.30-m reflector). The indicated correction to the prediction by S. Nakano in the 2006 Comet Handbook (and NK 946) is ΔT = +18.6 days.

2007 UT  α_{2000}  δ_{2000}  Mag. Observer
Jan. 9.07080 22^{h}53^{m}11.^{s}96 −5^{o}45′10′′.9 16.8 Kowalski
9.07407 22 53 12.66 −5 45 07.5 16.4  _{”}
9.07734 22 53 13.28 −5 45 04.2 16.8  _{”}
9.08061 22 53 13.78 −5 45 01.2 18.3  _{”}
9.09382 22 53 16.28 −5 44 45.3 17.0 Christensen
9.09470 22 53 16.42 −5 44 44.3  _{”}
9.09558 22 53 16.60 −5 44 43.5  _{”}
9.09647 22 53 16.74 −5 44 42.0  _{”}
10.07002 22 56 24.14 −5 25 43.1 17.3  _{”}
10.07782 22 56 25.67 −5 25 34.3  _{”}
10.08178 22 56 26.46 −5 25 29.2  _{”}
10.10260 22 56 30.40 −5 25 05.4 18.6 Hug
10.10699 22 56 31.20 −5 25 00.8 18.7  _{”}
10.11151 22 56 32.06 −5 24 56.5 19.1  _{”}
10.12584 22 56 34.79 −5 24 37.4 18.8 McGaha
10.13161 22 56 35.92 −5 24 30.3 19.1  _{”}
10.13738 22 56 37.10 −5 24 23.9 18.8  _{”}
2007 January 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green